# UVLM: Benchmarking Video Language Model for Underwater World Understanding

### v1

@@ Abstract
Recently, the remarkable success of large language models (LLMs) has achieved a profound impact on the field of artificial intelligence. Numerous advanced works based on LLMs have been proposed and applied in various scenarios. Among them, video language models (VidLMs) are particularly widely used. However, existing works primarily focus on terrestrial scenarios, overlooking the highly demanding application needs of underwater observation. To overcome this gap, we introduce UVLM, an under water observation benchmark which is build through a collaborative approach combining human expertise and AI models. To ensure data quality, we have conducted in-depth considerations from multiple perspectives. First, to address the unique challenges of underwater environments, we selected videos that represent typical underwater challenges including light variations, water turbidity, and diverse viewing angles to construct the dataset. Second, to ensure data diversity, the dataset covers a wide range of frame rates, resolutions, 419 classes of marine animals, and various static plants and terrains. Next, for task diversity, we adopted a structured design where observation targets are categorized into two major classes: biological and environmental. Each category includes content observation and change/action observation, totaling 20 distinct task types. Finally, we designed several challenging evaluation metrics to enable quantitative comparison and analysis of different methods. Experiments on two representative VidLMs demonstrate that fine-tuning VidLMs on UVLM significantly improves underwater world understanding while also showing potential for slight improvements on existing in-air VidLM benchmarks, such as VideoMME and Perception text. The dataset and prompt engineering will be released publicly.
1 Introduction
@@
Video-language understanding [1, 2] stands at the forefront of multimedia research, empowering systems to interpret, reason about, and generate natural language descriptions of temporal and dynamic visual content [3]. Recent advances [3, 4, 5] in video-language models (VidLMs) have demonstrated impressive performance in tasks such as video captioning, temporal grounding, and visual question answering, focusing primarily on human-centric scenarios and common object interactions. Despite these achievements, a critical question remains: Can current VidLMs effectively understand videos captured in special imaging conditions, such as underwater environments?
@@
This question is especially relevant as underwater environments constitute an uncharted domain with immense scientific value (e.g., marine biodiversity monitoring, ecosystem health assessment) and substantial engineering applications (e.g., autonomous underwater vehicles, offshore infrastructure inspection). As illustrated in Figure 1, applying VidLMs to underwater content presents three critical challenges that distinguish this domain from conventional video-language tasks:
@@
  • Degraded Visual Features Hindering Analytical Decisions. The underwater domain introduces fundamental perceptual barriers that confound existing video understanding approaches. Underwater scenes exhibit variable illumination with rapid light attenuation at depth, wavelength-dependent color distortion that shifts the visual spectrum, fluctuating turbidity affecting visibility [6, 7]. Standard VidLMs designed for terrestrial scenario environments struggle to perform effectively underwater due to these low-quality visual cues.
@@
  • Lack of Scientific Domain Knowledge. Underwater content requires specialized ecological expertise for accurate interpretation. Unlike common scenarios featuring familiar objects and actions, underwater videos capture complex species interactions, specialized behavioral patterns, and environmental relationships that demand expert knowledge to decode [8, 9]. The interpretation of underwater footage requires understanding multiple layers of information, including taxonomic identification, morphological characteristics, behavioral states, and environmental contexts. Models must capture the intricate ecological relationships between marine organisms’ appearance, behaviors, and habitats, creating a significant knowledge gap for systems trained primarily on common objects and human activities.
@@
  • Data Resource Scarcity. The development of effective video-language models for underwater environments is critically hindered by the absence of comprehensive training resources. Current video-language datasets predominantly focus on everyday human activities (HowTo100M [10]), sports (Sports-1M [11]), or general-purpose actions (Kinetics [12]), offering minimal support for specialized scientific domains. While underwater-specific datasets exist, they typically prioritize narrow tasks such as object tracking (WebUOT-1M [13]) or instance segmentation (UIIS [14]). These image-based datasets fail to equip models with the scientific domain knowledge necessary for comprehensive understanding of marine organism behaviors in underwater videos. This data gap creates a significant barrier to advancing video-language understanding beyond common scenarios to specialized scientific contexts.
@@
To bridge these gaps, a comprehensive Underwater Video-Language Multimodal (UVLM) dataset integrates professional annotations is presented. Compared to existing video language understanding benchmarks, our UVLM demonstrates distinct differences in both content composition and construction methodology. First, in terms of content, UVLM is the first video-language benchmark specifically designed for underwater environments. To ensure the dataset accurately captures the distinctive characteristics of underwater settings, we carefully selected videos that encompass the unique challenges of this domain, including low-light conditions, water turbidity, and the highly variable movement patterns of marine organisms. Second, regarding construction methodology, we adopted a structured framework combining human-AI collaboration. The annotation targets encompass both biological and environmental elements, covering static and dynamic scenarios. After frame-by-frame manual annotation, we leveraged these human-annotated labels to guide GPT-4o in generating diverse sample content. Finally, all data underwent manual verification, with inconsistent entries either re-annotated or supplemented using search engines for factual accuracy.
@@
The final dataset comprises 0.9M frames, 419 distinct biological categories, and diverse underwater scenarios, encompassing 20 types of video-language understanding tasks. Additionally, we established six specialized metrics specifically designed for quantitative performance comparison on the dataset. UVLM advances video-language research through its comprehensive coverage of core technical challenges:
@@
2 Related Work
Table 1: Representative video language benchmarks
Name Venue Themes & Scenarios
MSVD-QA [15], Next-QA [16] $M\prime$17, CVPR $\prime$21 Daily life
[…]
EgoSchema [18] NeurIPS $\prime$23 Life and human action, egocentric perspective
Video-Bench [19] arXiv $\prime$23 Television, music videos, NBA highlights
InternVid [20] ICLR $\prime$23 People and blogs, education, news and politics, how-to and style, entertainment, gaming, science, sports
MovieChat-1K [21] CVPR $\prime$24 Movie clips
MVBench [4] CVPR $\prime$24 Life, human action, movies, people, pets, science, instruction, finance, news, sports, humor, etc.
LONGVIDEOBENCH [22] NeurIPS $\prime$24 Life, movie, knowledge and news
E.T. Bench [23] NeurIPS $\prime$24 Indoor activities, tabletop, sports, egocentric views, cooking, news and vlogs, how-to, open-ended
LVBench [24] arXiv $\prime$24 Sports documentary, event record, lifestyle, TV show, cartoon
EgoThink [25] CVPR $\prime$24 Daily life understanding from first-person perspective
[…]
2.1 Video Language Understanding Benchmark
In recent years, many video language benchmarks have emerged, each targeting specific themes and application domains. Some focus primarily on everyday life scenarios  [15, 16, 17], while others emphasize human action or movie clips  [18, 21]. More comprehensive datasets cover a wider range of categories, including knowledge, sports, and instructional videos  [4, 22, 26, 27]. Table 1 provides a concise summary of representative benchmarks, highlighting their respective content coverage. Although these efforts have significantly advanced video-language research in terrestrial scenario contexts, the underwater environment remains largely uncharted, motivating our investigation of underwater video-language benchmark.
2.2 Underwater Observation Datasets
Low-level perception datasets focus on fundamental image enhancement and quality assessment, such as LSUI [28]. These datasets provide essential resources for developing and validating algorithms to address the typical degradation in underwater imagery. More recent efforts have extended this work into the temporal domain with video datasets like DRUVA [29], which contains 6,000 frames. Mid-level recognition tasks have also benefited from dedicated datasets. For instance, Wildfish [30] is tailored for marine life recognition, while COU [31] supports segmentation research. Additionally, a range of tracking datasets, such as VMAT [32] and WebUOT-1M [13] has further advanced object-centric underwater understanding. At the high-level, multimodal tasks are beginning to emerge. The MarineInst dataset [33] marks a critical step forward by supporting advanced tasks such as image segmentation and captioning, thereby opening new avenues for comprehensive underwater analysis.
Table 2: Comparison of underwater observation datasets.
Dataset Venue Im Vid Lang Seq Frame Task Categ.
@@ 2.2 Underwater Observation Datasets
[…]
UVLM - 2111 0.86M Video Understanding 419
2.3 Underwater Visual Analysis
Research efforts in this area [36] have focused primarily on three major directions: enhancement of underwater image quality [37, 38], interpretation of underwater objects [30, 31], estimation and modeling of the underwater environment [39, 40].
Underwater images typically have quality degradation caused by the scattering and absorption of light, color distortion, etc. To address these challenges, numerous underwater image enhancement algorithms [37, 41, 42, 38] have been proposed to restore visual information. Besides, for underwater perception, researchers have developed numerous specialized methods, including fish recognition [30], object detection [31], instance segmentation [43], and object tracking [13, 32], substantially advancing object-centric understanding in underwater environments.
Beyond the interpretation of objects, substantial progress has also been achieved in vision-based underwater environment estimation and modeling. Recent techniques such as dense 3D mapping [39], depth estimation [44, 45, 42] using diffusion models, and underwater scene reconstruction based on neural radiance fields [40] have greatly advanced our capabilities in comprehensive environmental understanding and reconstruction. Building upon the notable successes of single-modality underwater vision, Zheng et al. [33] further introduced multimodal approaches by incorporating textual information into underwater visual analysis.
In summary, underwater vision has seen steady progress, from low-level enhancement to high-level understanding. However, due to data limitation, the advanced VidLMs have yet to be effectively applied in underwater scenarios. We hope the proposed UVLM can help bridge this gap.
3 The proposed UVLM
@@ 3.1 Overview
In this paper, taking these challenges into consideration, we developed UVLM, the first underwater video-language benchmark. Firstly, we collect videos from typical underwater environments like oceans, lakes, and rivers, which exhibit different degradation problems such as water surface ripples, turbid water, and light scattering. Secondly, we select videos containing different content types, including static environmental features, static observations of organisms, dynamic scene variations, and dynamic biological behaviors, etc. To systematically inject domain knowledge into the dataset, we construct textual descriptions from 20 distinct perspectives covering observation content, behavioral analysis, and environmental dynamics. Thirdly, to maximize the scale of our dataset, we adopt a dual approach: systematically collecting data from online sources while strategically screening and re-annotating existing relevant datasets.
The final dataset comprises approximately two thousand carefully selected video sequences and 0.9 million frames, covering 419 different marine organisms and about 40 thousand video-language pairs. Previous underwater observation datasets [31, 43, 46, 47, 48, 49, 50, 51] typically focus on traditional computer vision tasks such as image segmentation and single object tracking, which are pure vision tasks. A few underwater observation datasets [33, 52] also incorporate visual and language elements, but the tasks are still limited to typical applications such as captioning. Unlike these datasets, the proposed UVLM is designed to incorporate specific underwater domain knowledge and requirements. To achieve this, we designed 16 to 20 relative questions referring to underwater research research topics such as marine organism recognition, behavioral analysis and prediction, habitat pattern characterization, subaquatic environmental monitoring, and ecosystem dynamics assessment, etc., covering both biological and environmental dimensions critical to underwater exploration and conservation. A comparison with previous underwater datasets is presented in Table 2.
3.2 Benchmark Construction
@@
The first path is collecting underwater videos from websites such as youtube and bilibili. We collect web-crawled data while enforcing two quality control principles during acquisition: 1) Both camera and observation objects or scenes must be under water. This selection rule is more in line with real world underwater vehicle and robot application scenarios; 2) The typical underwater scenarios and targets with different challenges and characteristics must be covered. The videos should cover the typical environments, like oceans, lakes, rivers and fish tanks, etc. Meanwhile, the chosen targets should include (but not limited to) common animals (fish, whales, prawns, tortoises, etc.), people, mermaids and generic objects. With the principles, we selected about 4 hundred video sequences, each of which contains 100 to 3000 frames, covering 53 categories of objects. The second path is re-annotating existing dataset. For scalability considerations, we selected WebUOT [13] as our primary data source.
@@
Manual annotation. As emphasized in our introduction, the substantial domain gap between terrestrial scenario and underwater environments requires auxiliary annotation information to effectively leverage existing VidLMs for collaborative labeling. To facilitate granular analysis of underwater targets, such as studying the behavioral traits and ecological patterns of marine organisms, we annotate each frame with both bounding boxes and fine-grained taxonomic classifications aligned with marine biology standards. These annotations function as priors when fed into AI models for text generation.
@@
For videos collected from websites, we employed 12 annotators annotate videos following three principles: 1) For occluded objects, only the visible part is marked with a rectangular box; 2) For the sharply protruding part, such as the tail and mouth of fish, whether to contain it in the bounding box is determined based on the proportion of the target and background in the additional introduced area. In specific, if the object’s protruding parts accounts for more than one-third of the additional introduced area, the sharply protruding part would be marked, otherwise it should not be marked; 3) Each frame is annotated with an axis aligned bounding box using the DarkLabel toolbox [^1] https://github.com/darkpgmr/DarkLabel, and each sequence has been assessed by three domain experts to minimize annotation errors. Unqualified labels would be sent to other annotators for re-labeling.
@@
For videos from WebUOT, we followed several steps to refine video quantity. We first performed data cleaning to filter out videos with information interference, such as those containing large areas of subtitles or watermarks that affect video quality. Then, we conducted scene cleaning to remove videos that were not in natural underwater environments, such as those filmed in aquariums, fish tanks, or simulated gaming scenes. Next, we adopted SAM [53] and LaMa [54] to remove small-area watermarks, subtitles, and other information interference. We use mouse clicks to select the area, SAM for segmentation, and then LaMa for inpainting. Next, to avoid challenges caused by excessively long videos and focus specifically on evaluating the impacts from underwater environments, we split some overly long videos into segments of 300-600 frames each. Finally, based on statistical analysis, we removed some videos to balance the distribution of different observed object categories. In post-processing, a total of 544 videos were filtered out.
@@
The questions include two formats: multiple-choice and open-ended. For instance, multiple-choice: "What marine species is primarily observed in the images? A) Clownfish B) Chromis dimidiatus C) Parrotfish". Open-ended: "What is the overall setting of the video, and how does it influence Chromis dimidiatus’ activities?". During the Q&A process, we implemented clear prompt constraints, including: 1) Topic-specific constraints (e.g., focusing on taxonomic details, observed behaviors, or physicochemical properties of the water body); 2) Style requirements (e.g., diversifying sentence structures and enriching question types). Each video generates 16 to 20 video-text pairs. Please refer to the Appendix A for our generation prompts.
@@
Stringent quality control was applied to all Q&A pairs generated by GPT-4o through a two-tier human review protocol. The first round, conducted by general reviewers, focused on detecting conflicts between the information supplied to GPT-4o and the content it produced. The second round involved senior experts who performed in-depth edits to ensure factual precision and domain conformity. During both rounds, hallucinations or statements irrelevant to the video context were either excised or fully rewritten. Whenever a model description diverged from the ground truth, we leveraged the accompanying assistant information to cross-validate and revise the erroneous content, thereby safeguarding the integrity of the dataset.
Figure 3: Statistics on UVLM. (a) Scene distribution; (b) Observation target distribution; (c) Object categories
3.3 UVLM Statistics
@@
3.4 Evaluation Metric

### v2

@@ Abstract
Recently, video-language models (VidLMs) have gained widespread attention and adoption. However, existing works primarily focus on terrestrial scenarios, overlooking the highly demanding application needs of underwater observation. To overcome this gap, we introduce UVLM, an under water observation benchmark which is build through a collaborative approach combining human expertise and AI models. To ensure data quality, we have conducted in-depth considerations from multiple perspectives. First, to address the unique challenges of underwater environments, we selected videos that represent typical underwater challenges including light variations, water turbidity, and diverse viewing angles to construct the dataset. Second, to ensure data diversity, the dataset covers a wide range of frame rates, resolutions, 419 classes of marine animals, and various static plants and terrains. Next, for task diversity, we adopted a structured design where observation targets are categorized into two major classes: biological and environmental. Each category includes content observation and change/action observation, totaling 20 subtask types. Finally, we designed several challenging evaluation metrics to enable quantitative comparison and analysis of different methods. Experiments on two representative VidLMs demonstrate that fine-tuning VidLMs on UVLM significantly improves underwater world understanding while also showing potential for slight improvements on existing in-air VidLM benchmarks, such as VideoMME and Perception text. The dataset and prompt are publicly available at: https://github.com/Cecilia-xue/UVLM-Benchmark.
Introduction
@@
Video-language understanding (Xu et al. 2016; Anne Hendricks et al. 2017) stands at the forefront of multimedia research, empowering systems to interpret, reason about, and generate natural language descriptions of temporal and dynamic visual content (Patraucean et al. 2023). Recent advances (Patraucean et al. 2023; Li et al. 2024; Zhang et al. 2025a) in video-language models (VidLMs) have demonstrated impressive performance in tasks such as video captioning, temporal grounding, and visual question answering, focusing primarily on human-centric scenarios and common object interactions. Despite these achievements, a critical question remains: Can current VidLMs effectively understand videos captured in special imaging conditions, such as underwater environments?
@@
This question is especially relevant as underwater environments constitute an uncharted domain with immense scientific value (e.g., marine biodiversity monitoring, ecosystem health assessment (Xue et al. 2024)) and substantial engineering applications (e.g., autonomous underwater vehicles, offshore infrastructure inspection). As illustrated in Figure 1, applying VidLMs to underwater content presents three challenges that distinguish this domain from conventional video-language tasks:
@@
  • Degraded Visual Features Hindering Analytical Decisions. The underwater domain introduces fundamental perceptual barriers that confound existing video understanding approaches. Underwater scenes exhibit variable illumination with rapid light attenuation at depth, wavelength-dependent color distortion that shifts the visual spectrum, fluctuating turbidity affecting visibility (Islam, Xia, and Sattar 2020; Akkaynak and Treibitz 2019). Standard VidLMs designed for terrestrial scenario environments struggle to perform effectively underwater due to these low-quality visual cues.
@@
  • Lack of Scientific Domain Knowledge. Underwater content requires specialized ecological expertise for accurate interpretation. Unlike common scenarios featuring familiar objects and actions, underwater videos capture complex species interactions, specialized behavioral patterns, and environmental relationships that demand expert knowledge to decode (Marks et al. 2022; Han et al. 2024). The interpretation of underwater footage requires understanding multiple layers of information, including taxonomic identification, morphological characteristics, behavioral states, and environmental contexts. Models must capture the intricate ecological relationships between marine organisms’ appearance, behaviors, and habitats, creating a significant knowledge gap for systems trained primarily on common objects and human activities.
@@
  • Data Resource Scarcity. The development of effective video-language models for underwater environments is critically hindered by the absence of comprehensive training resources. Current video-language datasets predominantly focus on everyday human activities (HowTo100M (Miech et al. 2019)), sports (Sports-1M (Karpathy et al. 2014)), or general-purpose actions (Kinetics (Carreira and Zisserman 2017)), offering minimal support for specialized scientific domains. While underwater-specific datasets exist, they typically prioritize narrow tasks such as object tracking (WebUOT (Zhang et al. 2024)) or instance segmentation (UIIS (Lian et al. 2023a)). These image-based datasets fail to equip models with scientific domain knowledge necessary for comprehensive understanding of marine organism behaviors in underwater videos. This data gap creates a major barrier to advancing video-language understanding beyond common scenarios to specialized scientific contexts.
@@
To bridge these gaps, a comprehensive Underwater Video-Language Multimodal (UVLM) dataset with professional annotations is presented. Compared to existing video language understanding benchmarks, UVLM demonstrates distinct differences in both content composition and construction methodology. First, in terms of content, UVLM is the first video-language benchmark specifically designed for underwater environments. To ensure the dataset accurately captures the distinctive characteristics of underwater settings, we carefully selected videos that encompass unique challenges of this domain, including low-light conditions, water turbidity, and the highly variable movement patterns of marine organisms. Second, regarding construction methodology, we adopted a structured framework combining human-AI collaboration. The annotation targets encompass both biological and environmental elements, covering static and dynamic scenarios. After frame-by-frame manual annotation, we leveraged these human-annotated labels to guide GPT-4o in generating diverse sample content. Finally, all data underwent manual verification, with inconsistent entries either re-annotated or supplemented using search engines for factual accuracy.
@@
The final dataset comprises 0.9M frames, 419 distinct biological categories, and diverse underwater scenarios, encompassing 20 types of video-language understanding tasks. Additionally, we established eight specialized metrics for quantitative performance comparison on the dataset. UVLM advances video-language research through its comprehensive coverage of core technical challenges:
@@
Dataset Venue Im Vid Lang OQA BioInf Seq Frame Task Categ
@@
UVLM - 2111 0.86M VU 419
Table 1: Comparison of recent underwater observation datasets. Im, Vid, Lang, OQA, BioInf, Seq, and Categ denote Image, Video, Language, Open-ended Question Answering, Taxonomic Classification Information in Biology, Sequence, and Category, respectively. In Task, IR, DE, OC, IS, SOT, CAP, SOD, and VU refer to Image Restoration, Depth Estimation, Object Counting, Instance Segmentation, Single Object Tracking, Captioning, Salient Object Detection, and Video Understanding, respectively.
Related Work
Video Language Understanding Benchmark
In recent years, many video language benchmarks have emerged, each targeting specific themes and application domains. Some focus primarily on everyday life scenarios  (Xu et al. 2017; Xiao et al. 2021; Yu et al. 2019), while others emphasize human action or movie clips  (Mangalam, Akshulakov, and Malik 2023; Song et al. 2024). More comprehensive datasets cover a wider range of categories, including knowledge, sports, and instructional videos  (Li et al. 2024; Wu et al. 2024; Fu et al. 2025; Wang et al. 2024). Table 1 in Appendix.B provides a concise summary of representative benchmarks, highlighting their respective content coverage. Although these efforts have significantly advanced video-language research in terrestrial scenario contexts, the underwater environment remains largely uncharted, motivating our investigation of underwater video-language benchmark.
Underwater Observation Datasets
Low-level perception datasets focus on fundamental image enhancement and quality assessment, such as LSUI (Peng et al. 2023). These datasets provide essential resources for developing and validating algorithms to address the typical degradation in underwater imagery. More recent efforts have extended this work into the temporal domain with video datasets like DRUVA (Varghese et al. 2023), which contains 6,000 frames. Mid-level recognition tasks have also benefited from dedicated datasets. For instance, Wildfish (Zhuang, Wang, and Qiao 2018) is tailored for marine life recognition, while COU (Mukherjee et al. 2025) supports segmentation research. Additionally, a range of tracking datasets, such as VMAT (Cai et al. 2023) and WebUOT (Zhang et al. 2024) has further advanced object-centric underwater understanding. At the high-level, multimodal tasks are beginning to emerge. The MarineInst dataset (Zheng et al. 2024) marks a critical step forward by supporting advanced tasks such as image segmentation and captioning, thereby opening new avenues for comprehensive underwater analysis.
@@
In this paper, taking these challenges into consideration, we developed UVLM, the first underwater video-language benchmark. Firstly, we collect videos from typical underwater environments like oceans, lakes, and rivers, which exhibit different degradation problems such as water surface ripples, turbid water, and light scattering. Secondly, we select videos containing different content types, including static environmental features, static observations of organisms, dynamic scene variations, and dynamic biological behaviors, etc. To systematically inject domain knowledge into the dataset, we design textual descriptions structured into 20 subtasks, reflecting VidLMs’ capabilities across 9 dimensions, such as marine animal behavior, water bodies, and geological features (details in Appendix.B). Thirdly, to maximize the scale of our dataset, we adopt a dual approach: systematically collecting data from online sources while strategically screening and re-annotating existing relevant datasets.
The final dataset comprises approximately two thousand carefully selected video sequences and 0.9 million frames, covering 419 different marine organisms and about 40 thousand video-language pairs. Previous underwater observation datasets (Lian et al. 2023b; Zhao et al. 2021; Alawode et al. 2022, 2023; Zhang et al. 2025b; Fan et al. 2019; Huang, Zhao, and Huang 2021) typically focus on traditional computer vision tasks such as image segmentation and single object tracking, which are pure vision tasks. A few underwater observation datasets (Zheng et al. 2024; Li et al. 2025) also incorporate visual and language elements, but the tasks are still limited to typical applications such as captioning. Unlike these datasets, the proposed UVLM is designed to incorporate specific underwater domain knowledge and requirements. To achieve this, we designed 16 to 20 relative questions referring to underwater research topics such as marine organism recognition, behavioral analysis and prediction, habitat pattern characterization, subaquatic environmental monitoring, and ecosystem dynamics assessment, etc., covering both biological and environmental dimensions critical to underwater exploration and conservation. A comparison with previous underwater datasets is shown in Table 1.
Figure 3: Statistics on UVLM. (a) Scene distribution; (b) Observation target distribution; (c) Several samples in UVLM; (d) Fine-grained taxonomic classification information (partial categories), from left to right: Kingdom, Phylum, Class, Order.
Data Collection and Annotation
@@
The first path is collecting underwater videos from websites such as youtube and bilibili. We collect web-crawled data while enforcing two quality control principles during acquisition: 1) Both camera and observation objects or scenes must be under water. This selection rule is more aligned with real world underwater vehicle and robot application scenarios; 2) The typical underwater scenarios and targets with different challenges and characteristics must be covered. The videos should cover typical environments, like oceans, lakes, rivers and fish tanks, etc. Meanwhile, the chosen targets should include (but not limited to) common animals (fish, whales, prawns, tortoises, etc.), people, mermaids and generic objects. With the principles, we selected about 400 video sequences, each of which contains 100 to 3000 frames, covering 53 classes of objects. The second path is re-annotating existing dataset. For scalability considerations, we selected WebUOT (Zhang et al. 2024) as our primary data source.
@@
Manual annotation. As emphasized in the Introduction, the substantial domain gap between terrestrial and underwater environments requires auxiliary annotation information to effectively leverage existing VidLMs for collaborative labeling. To facilitate granular analysis of underwater targets, such as studying the behavioral traits and ecological patterns of marine organisms, we annotate each frame with both bounding boxes and fine-grained taxonomic classifications aligned with marine biology standards. These annotations function as priors when fed into AI models for text generation.
@@
For videos collected from websites, we employed 12 annotators annotate videos following three principles: 1) For occluded objects, only the visible part is marked with a rectangular box; 2) For the sharply protruding part, such as the tail and mouth of fish, whether to contain it in the bounding box is determined based on the proportion of the target and background in the additional introduced area. In specific, if the object’s protruding parts accounts for more than one-third of the additional introduced area, the sharply protruding part would be marked, otherwise it should not be marked; 3) Each frame is annotated with an axis aligned bounding box using the DarkLabel toolbox [^1] https://github.com/darkpgmr/DarkLabel. Each sequence is assessed by three domain experts to minimize annotation errors. Unqualified labels would be sent to other annotators for re-labeling.
@@
For videos from WebUOT, we followed several steps to refine video quantity. We first performed data cleaning to filter out videos with information interference, such as those containing large areas of subtitles or watermarks that affect video quality. Then, we conducted scene cleaning to remove videos that were not in natural underwater environments, such as those filmed in aquariums, fish tanks, or simulated gaming scenes. Next, we adopted SAM (Kirillov et al. 2023) and LaMa (Suvorov et al. 2022) to remove small-area watermarks, subtitles, and other information interference. We use mouse clicks to select the area, SAM for segmentation, and then LaMa for inpainting. Next, to avoid challenges caused by excessively long videos and focus specifically on evaluating the impacts from underwater environments, We segmented some longer videos with fewer observed organisms into 300-600 frame clips to balance the sample distribution. Some videos collected from the internet even contain up to 3,000 frames, so the dataset can also evaluate the models’ ability to understand extended underwater scenarios.
To achieve reliable and fine-grained taxonomic classification of marine animals in these videos, we implemented a structured three-phase annotation procedure. In the first phase, four annotators with extensive marine biology knowledge independently labeled each target within the video frames, providing both species-level identifications and detailed taxonomic classifications according to the classic five-kingdom system (kingdom, phylum, class, order, etc.) in biology (Whittaker 1969), supported by authoritative sources such as Wikipedia. In the second phase, annotations were cross-validated by annotator pairs. Any disagreements were resolved by consulting a third annotator for majority consensus, with unresolved cases flagged for expert review. In the third phase, a senior marine biology expert reviewed the validated annotations, marking questionable cases. Finally, all flagged annotations were collectively discussed by five experts to establish definitive classifications. This rigorous approach ensured high accuracy and consistency in fine-grained taxonomic annotation.
Method Objective Metrics LLM-based Judgement Metrics
MCA FGC SA DC VPA EDA SBM Overall Accuracy
Closed-source VidLMs and other Open Source Large VidLMs
GPT-4o 77.72 81.47 77.67 73.40 78.23 80.07 79.73 77.95
Claude3.7-Sonnet 76.61 82.64 73.35 73.58 74.10 79.71 76.35 76.09
Gemini2.5-Flash 78.22 86.27 72.43 73.34 70.53 78.32 74.92 75.00
Qwen2.5VL-72B 75.97 80.57 74.22 71.94 74.85 78.45 77.40 75.49
Base VLM
Qvis2.5-2B 61.87 29.85 52.88 49.70 54.66 55.81 50.62 53.06
Qvis2.5-2B + UVLM 72.89 59.25 67.64 62.39 65.30 69.05 66.29 68.85 (+15.79)
Base VidLMs
InternVL2.5-1B 48.54 29.61 44.35 47.30 45.30 50.15 44.85 46.73
VideoLLaMA3-2B 57.17 31.25 55.89 58.48 58.53 63.15 55.74 58.39
Qwen2.5VL-2B 59.47 35.23 54.12 50.65 53.47 57.31 50.60 52.97
InternVL2.5-8B 57.64 36.63 57.70 59.75 59.45 63.48 61.05 60.15
VideoLLaMA3-7B 63.83 42.62 60.43 62.07 62.05 66.97 61.41 62.70
Qwen2.5VL-7B 66.22 48.36 66.67 59.98 61.74 68.41 61.79 63.57
InternVL2.5-1B + UVLM 64.52 45.37 54.48 56.55 55.78 65.16 58.74 59.14 (+12.41)
VideoLLaMA3-2B + UVLM 70.41 46.35 63.54 64.76 66.71 70.28 67.32 66.67 (+8.28)
Qwen2.5VL-2B + UVLM 62.38 56.08 60.47 55.81 57.93 61.45 58.26 58.44 (+5.47)
InternVL2.5-8B + UVLM 70.26 43.94 65.66 66.42 65.05 71.10 69.98 69.45 (+9.30)
VideoLLaMA3-7B + UVLM 76.85 57.17 70.88 70.17 70.40 76.35 73.66 73.04 (+10.34)
Qwen2.5VL-7B + UVLM 71.69 63.41 70.47 67.25 67.29 72.16 65.76 68.08 (+4.51)
Table 2: Performance comparison of different methods on UVLM test set. Metric abbreviations: MCA, FGC, SA, DC, VPA, EDA, SBM denote Multiple Choice Accuracy, Fine-grained Taxonomic Classification, Semantic Accuracy, Detail Completeness, Visual Perception Accuracy, Environmental Description Accuracy, and Species Behavior Matching, respectively.
@@
The questions include two formats: multiple-choice and open-ended. For instance, multiple-choice: ”What marine species is primarily observed in the images? A) Clownfish B) Chromis dimidiatus C) Parrotfish”. Open-ended: ”What is the overall setting of the video, and how does it influence Chromis dimidiatus’ activities?”. During the Q&A process, we implemented clear prompt constraints, including: 1) Topic-specific constraints (e.g., focusing on taxonomic details, observed behaviors, or physicochemical properties of the water body); 2) Style requirements (e.g., diversifying sentence structures and enriching question types). Each video generates 16 to 20 video-text pairs. Appendix.A offers the generation prompts.
@@
Stringent quality control was applied to all Q&A pairs generated by GPT-4o through a two-tier human review protocol. The first round, conducted by general reviewers, focused on detecting conflicts between the information supplied to GPT-4o and the content it produced. The second round involved senior experts performing in-depth edits to ensure factual precision and domain conformity. During both rounds, hallucinations or statements irrelevant to the video context were either removed or rewritten. Whenever a model description diverged from ground truth, we leveraged the accompanying assistant information to cross-validate and revise the erroneous content, safeguarding dataset integrity.
UVLM Statistics
@@
Figure 4: Visualization examples of fine-tuning with UVLM. Error and correct descriptions are marked in red and green.
Evaluation Metrics